1. Solve the ordinary linear differential equation by the method of undetermined coefficients: y'' -2y' -3y=2e^ 4x


1
Expert's answer
2022-03-23T08:28:21-0400

Solution

For the homogeneous equation y'' - 2y' - 3y = 0 the characteristic equation is

λ2 - 2λ - 3=0  => "\lambda_{1,2}=1\pm\sqrt{1+3}"   => λ1 = -1, λ2 = 3

So the solution of homogeneous equation is y0(x) = C1e-x+C2e3x, where C1, C2 are arbitrary constants.

Partial solution may be find in the form y1(x) = Ae4x. Substitution into equation y'' -2y' -3y=2e4x gives

(16Ae4x -2*4A-3A) e4x  = 2e4x => 5A = 2 => A = 0.4   

Therefore the solution of given equation is

y(x) = y1(x)+y2(x) = C1e-x+C2e3x +0.4 e4x